OpenStudy (anonymous):

An underground mine elevator descends 12% of the total depth every minute. The total depth of the elevator is 75 meters below the ground. How many meters below the ground is the elevator after 2 minutes? A: -18 meters B: -9 meters C: -4.5 meters D: -22.5 meters

OpenStudy (anonymous):

@vera_ewing can you please help me :) Please?

OpenStudy (mathstudent55):

If it descends 12% of the total length in 1 minute, how much percent does it descend in 2 minutes?

OpenStudy (anonymous):

24% @mathstudent55

OpenStudy (mathstudent55):

Correct. Now you need to find what 24% of 75 meters is.

OpenStudy (mathstudent55):

What is 24% of 75 = ? To find a percent of a number, divide the percent by 100 and multiply by the number. 24% of 75 = = 24/100 * 75 = 0.24 * 75

OpenStudy (anonymous):

which is 18?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

so its -18?

jimthompson5910 (jim_thompson5910):

Correct. It descends 9 m every min, so after 2 min, it's at the -18 m mark.

OpenStudy (anonymous):

Thank you Guys Soooooo much I r eally Apreciate it :) @mathstudent55 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (mathstudent55):

Yes, since 24% of 75 m is 18 m, it descends 18 m in the first two minutes. Since it is descending, it is losing altitude, so you can call the displacement -18 m. (Strictly speaking, the way the question was asked, the answer should be 18 m, not -18 m, but you don't have such a choice.)

OpenStudy (mathstudent55):

You're welcome.